### v1

# Ensemble YOLO Framework for Multi-Domain Mitotic Figure Detection in Histopathology Images

Abstract
Accurate detection of mitotic figures in whole slide histopathological images remains a challenging task due to their scarcity, morphological heterogeneity, and the variability introduced by tissue preparation and staining protocols. The MIDOG competition series provides standardized benchmarks for evaluating detection approaches across diverse domains, thus motivating the development of generalizable deep learning models. In this work, we investigate the performance of two modern one-stage detectors, YOLOv5 and YOLOv8, trained on MIDOG++, CMC, and CCMCT datasets. To enhance robustness, training incorporated stain-invariant color perturbations and texture-preserving augmentations. In internal validation, YOLOv5 achieved superior precision, while YOLOv8 provided improved recall, reflecting architectural trade-offs between anchor-based and anchor-free detection. To capitalize on these complementary strengths, we employed an ensemble of the two models, which improved sensitivity without a major reduction in precision. These findings highlight the effectiveness of ensemble strategies built upon contemporary object detectors to advance automated mitosis detection in digital pathology.
keywords:
Mitosis Detection | Digital Pathology | YOLOv5 | YOLOv8 | Ensemble Learning | MIDOG 2025
{corrauthor}
navyasri.kelam@airamatrix.com
Introduction
Accurate mitosis detection is essential for histopathological diagnosis, directly impacting tumor classification and treatment planning. Challenges arise from the rarity of mitotic figures, their resemblance to non-mitotic structures, and staining variability across labs. The MICCAI MIDOG challenge (1) was designed to promote algorithms that generalize across domains with high accuracy. Automated mitosis detection has advanced rapidly, largely due to the MIDOG challenges, which benchmark cross-domain performance. MIDOG 2021 focused on breast cancer, enabling domain-specific generalization but limiting robustness. MIDOG++ (2) and MIDOG 2025 (3) introduced multi-domain datasets with diverse tissues, scanners, and staining protocols, making generalization a central challenge. Deep learning-based object detectors, especially the YOLO family, have become standard for mitosis detection due to their efficiency and accuracy. This study compares YOLOv5 and YOLOv8 on a curated multi-center dataset, examining how augmentation, dataset design, and architecture affect precision–recall dynamics, with a focus on model generalizability in heterogeneous tissue settings.
Related Work
The MIDOG challenge series has been pivotal in benchmarking mitosis detection algorithms under domain shift conditions. The MIDOG 2021 challenge (4) focused exclusively on breast cancer tissue and highlighted that domain generalization—arising from variability in tissue preparation, laboratory protocols, and scanner technologies—remains a central challenge in histopathological mitosis detection. While participants achieved high F1 scores, performance deteriorated on images from unseen scanners or conditions. Building on this, the MIDOG 2022 challenge (1) expanded to include multiple tumor types, allowing the evaluation of generalization strategies across more diverse histopathological domains. Participants were allowed to use publicly available mitosis datasets in addition to official training data, but the report noted that domain shifts continued to challenge algorithmic performance. Further extending these efforts, the MIDOG++ dataset and benchmark (2) incorporated a wider variety of tumor types and scanner modalities, and investigated training strategies such as leave-one-domain-out experiments to systematically assess domain generalization. These studies collectively underscore that increasing the diversity of training data is crucial for achieving robust and transferable mitosis detection performance across heterogeneous histopathological settings.
Datasets and Preprocessing
This study leverages three mitosis detection datasets to evaluate domain generalization and cross-species robustness. The MIDOG25 dataset(3) spans seven domains, including human and canine tumors, providing a diverse multi-scanner, multi-tumor benchmark for generalization. The MITOS_WSI_CMC dataset(5), focused on canine mammary carcinoma, introduces a cross-species domain shift. Meanwhile, the MITOS_WSI_CCMCT dataset(6), comprising canine cutaneous mast cell tumor slides, adds further complexity through interspecies and tumor-type variability. Following the MIDOG25 challenge protocol, the MIDOG25 dataset was first divided into training and testing subsets, with 80% of ROIs allocated for training and 20% reserved for evaluation. For the CMC and CCMCT datasets, three to five whole-slide images per center were selectively chosen to ensure histological heterogeneity, including tumor-dense regions, fibrotic tissue, necrotic zones, and normal stroma.
Augmentation
Augmentation was pivotal in this study, given the sensitivity of deep learning models in digital pathology to staining and imaging variability. Drawing on Litjens et al. (2017) (7), we implemented a diverse set of augmentations that simulate realistic conditions while maintaining the morphological integrity of mitotic figures. To address staining heterogeneity across labs, we applied color augmentations specifically hue, saturation, and brightness shifts to mimic variations in hematoxylin and eosin concentrations. These transformations help models prioritize structural features over raw color intensity.Texture-based augmentations were introduced to reflect acquisition-level inconsistencies. Gaussian blur simulated out-of-focus regions caused by tissue thickness or scanner optics, while sharpening enhanced nuclear edges to aid in distinguishing mitotic chromatin from apoptotic or necrotic nuclei. Gaussian noise was added to replicate scanner noise and preparation artifacts. We also incorporated advanced compositional augmentations. Mosaic augmentation (8) combined four random images into a single composite, enabling the model to learn across varied scales and contexts especially useful for detecting rare events like mitoses. Cutmix (9) exposed the model to partially visible nuclei, improving robustness in cases where mitotic figures are only partially captured. Together, these augmentations enhanced the model’s ability to generalize across domains, making it more resilient to the diverse challenges encountered in real world histopathology.
Model Architectures
YOLOv5-l(10) employs a CSPDarknet(11) backbone with an anchor-based detection head. The network is optimized for efficiency and precision, excelling in scenarios where false positives must be minimized. However, its reliance on predefined anchors can occasionally hinder performance on objects with highly variable scales, such as mitotic figures.
YOLOv8-m(12) introduces several critical improvements over its predecessor. Most notably, it adopts an anchor-free detection paradigm with a decoupled head, separating classification and localization tasks. This design not only reduces computational overhead, but also improves the ability of the model to recall rare and small objects. Enhanced feature aggregation within YOLOv8-m further facilitates gradient propagation, which is particularly beneficial for fine-grained mitotic detection. Collectively, these architectural refinements enable YOLOv8-m to achieve higher recall, although at the cost of slightly reduced precision compared to YOLOv5-l.
Model Training and Ensembling
Both YOLOv5-l and YOLOv8-m were trained independently for 200 epochs with an input resolution of 1024 × 1024. The Adam optimizer was employed with an initial learning rate of 0.01, decayed using a cosine annealing scheduler to stabilize convergence. Loss functions included bounding box regression, classification loss, and distribution focal loss. To further improve training stability and model performance, hyperparameters such as learning rate, augmentation probabilities, and loss coefficients were optimized using the Ultralytics Tune module (13). This automated tuning procedure enabled a systematic exploration of the hyperparameter space, ensuring that both models were trained under near-optimal conditions.
The choice of YOLOv5 and YOLOv8 was motivated by their architectural differences, which yield complementary detection behaviors. YOLOv5 which employs a CSPDarknet backbone with a PANet neck and an anchor-based detection head, optimized with CIoU and binary cross-entropy losses. YOLOv8, in contrast, introduces a CSP-Next backbone(14) with a decoupled head, adopts an anchor-free detection paradigm, and leverages advanced objectives such as the task aligned assigner and the distribution focal loss. These differences result in distinct inductive biases: YOLOv5 tends to favor well-structured mitotic figures that conform to anchor priors, producing more conservative predictions with higher precision, whereas YOLOv8 exhibits greater flexibility in localizing smaller or irregular mitoses, often leading to higher recall.
Model performance was assessed using precision, recall and the F1 score, as these jointly capture the trade-off between false positives and false negatives in the highly imbalanced mitosis detection task. The results showed that YOLOv5 provided greater precision, while YOLOv8 offered improved recall, confirming their complementary strengths. To exploit this complementarity, the final submission employed an ensemble of YOLOv5 and YOLOv8, combining their predictions to achieve higher sensitivity without substantially sacrificing precision.
Results and Discussion
In internal validation, YOLOv5 achieved a precision of 84. 314%, a recall of 79. 277%, and an F1 score of 81.711, reflecting its conservative detection behavior that reduces false positives but misses some true mitotic figures. In contrast, YOLOv8 obtained a precision of 82. 869%, a recall of 82. 610%, and an F1 score of 82.739, indicating a more balanced trade-off, but with slightly lower precision. These differences can be traced to their architectural designs: YOLOv5 employs a CSPDarknet53 backbone with a PANet neck and an anchor-based detection head, which favors well-structured mitotic figures that conform to anchor priors. YOLOv8, by comparison, introduces a CSP-Next backbone with a decoupled head, adopts an anchor-free detection paradigm, and incorporates advanced objectives such as the Task-Aligned Assigner and the Distribution Focal Loss. This design gives greater flexibility in localizing smaller or irregular mitoses, which often leads to higher recall. In combination, these complementary behaviors suggest that YOLOv5 tends to capture consistent, clearly delineated mitoses, while YOLOv8 recovers the true positives that YOLOv5 overlooks. To take advantage of this complementarity, an ensemble of the two models was constructed, resulting in a precision of 81. 107%, a recall of 85. 248%, and an F1 score of 83.126. The ensemble thus yielded improved sensitivity while maintaining competitive precision, underscoring the advantage of integrating detectors with complementary strengths in the mitosis detection task.
Acknowledgements
The authors acknowledge the organizers of the MIDOG25 challenge for providing a well-structured platform to advance research in medical imaging diagnostics. Their efforts in acquiring high-quality datasets and fostering collaborative innovation have been instrumental to this work.
We also extend our sincere thanks to my organization Airamatrix pvt ltd for supporting and encouraging our participation in this challenge. The opportunity to contribute to this initiative reflects the organization’s commitment to driving impactful solutions in AI-powered healthcare.